OpenStudy (anonymous):

Betsy is analyzing a quadratic function f(x) and a linear function g(x). Will they intersect? f(x) g(x) graph of the function f of x equals one third times x squared, minus 2 x g(x) 1 1 2 2 3 3 Yes, at positive x-coordinates Yes, at negative x-coordinates Yes, at negative and positive x-coordinates No, they will not intersect @Directrix

OpenStudy (solomonzelman):

f(x) = (1/3)x^2-2 and can you figure what the g(x) is based on the table you have?

OpenStudy (solomonzelman):

(x is equal to y in every point)

OpenStudy (solomonzelman):

you have points (1,1) , (2,2) and (3,3) (and will have (4,4) (5,5) and on) there you know that the x is equal to y, every time. so the line is just y=x.

OpenStudy (solomonzelman):

so, g(x)=x

OpenStudy (solomonzelman):

will they intersect? Set a system of equation: y = (1/3)x^2-2 y=x

OpenStudy (solomonzelman):

then solve for x, using substitution.

OpenStudy (anonymous):

is the answer B then? @SolomonZelman

OpenStudy (solomonzelman):

did you solve the system? how?

OpenStudy (anonymous):

I plotted the points and got B

OpenStudy (solomonzelman):

and I didn't get B.

OpenStudy (solomonzelman):

please solve the system/